OpenStudy (anonymous):

@iGreen 2x+6y=32 5x-2y=12 solve using substitution

OpenStudy (igreen):

2x + 6y = 32 Subtract 6y to both sides: 2x = -6y + 32 Divide 2 to both sides: x = -3y + 16 Plug in '-3y + 16' for 'x' in the 2nd equation: 5x - 2y = 12 5(-3y + 16) - 2y = 12 Distribute 5 into the parenthesis: -15y + 80 - 2y = 12 Combine like terms: -17y + 80 = 12 Subtract 80 to both sides, what's 12 - 80?

OpenStudy (anonymous):

-68?

OpenStudy (igreen):

Yes, so we have: -17y = -68 Divide -17 to both sides, what's -68 / -17?

OpenStudy (anonymous):

is the final answer (4,4)?

OpenStudy (igreen):

Yep, you got it.. (4, 4) is the final answer.

OpenStudy (anonymous):

okay I have two more problems I need help with. can u help?